OpenStudy (anonymous):

m=-3/2 passes through (-2,-2)

OpenStudy (jdoe0001):

ok, thanks for sharing that

OpenStudy (anonymous):

put it in point slope form \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

still don't get it

OpenStudy (anonymous):

substitute in the numbers the coordinates are expressed as (x1,y1) x1=-2 y1=-2

OpenStudy (anonymous):

y = - x - 5

OpenStudy (anonymous):

y=-3/2x-5